OpenStudy (anonymous):

Find the sum of the infinite geometric series 5+ 5/3 + 5/9 + 5/27 + if it exists.

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

5/(1-1/3)

OpenStudy (anonymous):

i got 7.46 or 500/67

OpenStudy (welshfella):

this is an other GS with r = 1/3

OpenStudy (welshfella):

5 / (2/3) = 5*3/ 2 = 7.5

OpenStudy (welshfella):

again this series converges to 7.5 so the sum to infinity does exist

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

yw but i must go now sorry

OpenStudy (anonymous):

okay